CVP Analysis and Special Decisions
Sweet Grove Citrus Company buys a variety of citrus fruit from growers and then processes the fruit into a product line of fresh fruit, juices, and fruit flavorings. The most recent year's sales revenue was $4,200,000. Variable costs were 60 percent of sales and fixed costs totaled $1,300,000. Sweet Grove is evaluating two alternatives designed to enhance profitability.

  • One staff member has proposed that Sweet Grove purchase more automated processing equipment. This strategy would increase fixed costs by $300,000 but decrease variable costs to 54 percent of sales.
  • Another staff member has suggested that Sweet Grove rely more on outsourcing for fruit processing. This would reduce fixed costs by $300,000 but increase variable costs to 65 percent of sales.

Round your answers to the nearest whole number.

(a) What is the current break-even point in sales dollars?
$Answer



(b) Assuming an income tax rate of 34 percent, what dollar sales volume is currently required to obtain an after-tax profit of $500,000?
$Answer

(c) In the absence of income taxes, at what sales volume will both alternatives (automation and outsourcing) provide the same profit?
$Answer

(d) Briefly describe one strength and one weakness of both the automation and the outsourcing alternatives.

Automation has less risk and a lower break-even point.

Outsourcing has higher profits if sales increase.

Automation has higher profits if sales increase. Outsourcing has less risk and a lower break-even point.

Automation has less risk. Outsourcing has higher profits if sales increase and a lower break-even point. Automation has higher profits if sales increase and a lower break-even point. Outsourcing has less risk.
Answer #1

a.Break even point in dollar sales = Fixed costs/Contribution Margin Ratio

= 1,300,000/40%

= $3,250,000

b.Desired Profit before tax = 500,000/(1-34%) = $757,575.76

Fixed costs = $1,300,000

Desired Contribution Margin = $2,057,575.76

Dollar sales volume required = 2,057,575.76/40%

= $5,143,939.39

c.Let the volume be x

x*40% - 1,300,000 = x*46% - 1,600,000

x = 5,000,000

d. Automation has higher profits if sales increase. Outsourcing has less risk and a lower break-even point.